OpenStudy (anonymous):

What is the common difference in the sequence 50, 40, 30, 20, 10, …

OpenStudy (anonymous):

10

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

-10..

OpenStudy (anonymous):

actually you doint have to put -10 because the question already said common difference

OpenStudy (anonymous):

Then what about 10 ,20 ,30 ,40 @severus4ever

OpenStudy (anonymous):

@╰☆╮Openstudier╰☆╮ , i dont get your question, then what about it?

OpenStudy (anonymous):

It would be +10 in my question but in the question she asked it would be -10 as it is reducing by 10

OpenStudy (anonymous):

-10 *

OpenStudy (anonymous):

yes i agree with that, but you dont neccessarily have to put -10 if it already said whats the common DIFFERENCE, which automatically puts a negative sign next to the anser. If you put -10, then the answer would have been turning it into a positive making it increase rather than decrease.

OpenStudy (anonymous):

thats what i was confused about ... if it is multiple choice and one option is 10 and another is -10 then which would I choose .

OpenStudy (anonymous):

Check which one is correct. trial and error

OpenStudy (tkhunny):

You choose the correct one. It is convention to proceed left to right. 50, 40, 30, 20, 10, 40-50 = -10 30-40 = -10 20-30 = -10 10-20 = -10

OpenStudy (anonymous):

its -10 not 10 ...

OpenStudy (anonymous):

did you submit it and see which one is right?

OpenStudy (tkhunny):

If -10 is not accepted, please dispute the homework assignment. There had better be a very good explanation for abandoning convention in this matter. I would be happy to argue it. Perhaps you live in a culture where it is common to read right to left? That might create a different outcome.

OpenStudy (anonymous):

@tkhunny, was that second comment directing towards me?

OpenStudy (tkhunny):

It is an odd thing, culture. We all tend to believe that our present culture is the only one that exists. It is important to be sensitive to other cultures. I'm just thinking on how +10 could be correct. I thought of one way - a different culture.

OpenStudy (anonymous):

im not saying aznything about +10, the answer already said common DIFFERENCE so you wouldn't have to put a negative sign if it already is there, difference = -